# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 2.4. Truncating $\infty$-operads
For every ${\mathrm{n}}\geq 1$ let $\Sigma_{\geq 1}^{\leq{\mathrm{n}}}\subset\Sigma_{\geq 1}$ be the full subcategory of sets with less or equal than ${\mathrm{n}}$ elements.
The embedding $\iota:\Sigma_{\geq 1}^{\leq{\mathrm{n}}}\subset\Sigma_{\geq 1}$ induces a localization
$\iota^{\ast}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\rightleftarrows{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{\mathrm{n}}}:\iota_{\ast},$
where the left adjoint is restriction and the fully faithful right adjoint extends beyond degree ${\mathrm{n}}$ by the zero object.
@@
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category compatible with the zero object that admits small colimits. For every ${\mathrm{n}}\geq 1$ the localization $\iota^{\ast}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\rightleftarrows{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{\mathrm{n}}}:\iota_{\ast}$ is compatible with the composition product on ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$
@@
By Proposition 7.9 we find that ${\mathcal{C}}$ embeds symmetric monoidally into a symmetric monoidal $\infty$-category ${\mathcal{C}}^{\prime}$ compatible with small colimits such that the embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ preserves the zero object and admits a left adjoint ${\mathrm{L}}.$ We obtain an embedding ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}_{\geq 1}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}_{\geq 1}$ of non-symmetric $\infty$-operads and an embedding ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{\mathrm{n}}}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}^{\leq{\mathrm{n}}}.$ Thus ${\mathrm{X}}_{1}\circ...\circ{\mathrm{X}}_{\mathrm{k}}$ is the image of the corresponding composition product in ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}$ under the functor ${\mathrm{L}}_{\ast}:{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ induced by ${\mathrm{L}}.$ A morphism of ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ is a local equivalence if and only if it induces an equivalence in degrees $\leq{\mathrm{n}}$ and similar for ${\mathcal{C}}^{\prime}.$ So the embedding ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}$ and its left adjoint ${\mathrm{L}}_{\ast}:{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ preserve local equivalences. Hence we can reduce to the case that the symmetric monoidal structure on ${\mathcal{C}}$ is compatible with small colimits.
Let $\mathsf{f}:{\mathrm{X}}\to{\mathrm{Y}}$ be a morphism in ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ such that $\mathsf{f}_{\mathrm{j}}:{\mathrm{X}}_{\mathrm{j}}\to{\mathrm{Y}}_{\mathrm{j}}$ is an equivalence for every ${\mathrm{j}}\leq{\mathrm{n}}$ and ${\mathrm{Z}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$ We want to see that $(\mathsf{f}\circ{\mathrm{Z}})_{\mathsf{s}}$ and $({\mathrm{Z}}\circ\mathsf{f})_{\mathsf{s}}$ are both equivalences for every $\mathsf{s}\leq{\mathrm{n}}.$ We compute
$(\mathsf{f}\circ{\mathrm{Z}})_{\mathsf{s}}=\underset{{\mathrm{k}}\geq 0}{\coprod}(\underset{{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}=\mathsf{s}}}{\coprod}\mathsf{f}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Z}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}=\underset{\mathsf{s}\geq{\mathrm{k}}\geq 0}{\coprod}(\underset{{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}=\mathsf{s}}}{\coprod}\mathsf{f}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Z}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}},$
$({\mathrm{Z}}\circ\mathsf{f})_{\mathsf{s}}=\underset{{\mathrm{k}}\geq 0}{\coprod}(\underset{{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}=\mathsf{s}}}{\coprod}{\mathrm{Z}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}\mathsf{f}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}.$
@@
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category that admits small colimits and a zero object preserved by the tensor product in each component. For every ${\mathrm{n}}\geq 1$ let ${\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq{\mathrm{n}}}\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}$ be the full subcategory of $\infty$-operads in ${\mathcal{C}}$ whose underlying symmetric sequence vanishes beyond degree ${\mathrm{n}},$ i.e. belongs to the full subcategory ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{\mathrm{n}}}\subset{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$
@@
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category that admits small colimits and a zero object preserved by the tensor product in each component. The embedding ${\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq{\mathrm{n}}}\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}$ admits a left adjoint $(-)_{\leq{\mathrm{n}}}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq{\mathrm{n}}}$ that fits into a commutative square, where the bottom functor is restriction:
${{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq{\mathrm{n}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{\mathrm{n}}}.}$
@@
If ${\mathcal{C}}$ is a presentably symmetric monoidal $\infty$-category, by the adjoint functor theorem the functor $(-)_{\leq{\mathrm{n}}}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq{\mathrm{n}}}$ preserving small limits and sifted colimits admits a left adjoint denoted by $\mathsf{f}_{\mathrm{n}}$. The composition
$\mathsf{f}_{\mathrm{n}}\circ(-)_{\leq{\mathrm{n}}}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq{\mathrm{n}}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}$
is left adjoint to the functor $\tau_{\mathrm{n}}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\xrightarrow{(-)_{\leq{\mathrm{n}}}}{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq{\mathrm{n}}}\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}.$
@@
$\mathsf{f}_{1}({\mathcal{O}}_{\leq 1})\to...\to\mathsf{f}_{\mathrm{n}}({\mathcal{O}}_{\leq{\mathrm{n}}})\to\mathsf{f}_{{\mathrm{n}}+1}({\mathcal{O}}_{\leq{\mathrm{n}}+1})\to...\to{\mathcal{O}}$
is a colimit diagram (see [21, Remark 4.23.] for details).
@@
${\mathcal{O}}\circ_{\mathsf{f}_{1}({\mathcal{O}}_{\leq 1})}{\mathrm{A}}\to...\to{\mathcal{O}}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{O}}_{\leq{\mathrm{n}}})}{\mathrm{A}}\to{\mathcal{O}}\circ_{\mathsf{f}_{{\mathrm{n}}+1}({\mathcal{O}}_{\leq{\mathrm{n}}+1})}{\mathrm{A}}\to...\to{\mathcal{O}}\circ_{\mathcal{O}}{\mathrm{A}}\simeq{\mathrm{A}}$
@@
$\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})({\mathrm{A}},{\mathrm{B}})\to\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})(\mathrm{colim}_{{\mathrm{n}}\geq 1}{\mathcal{O}}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{O}}_{\leq{\mathrm{n}}})}{\mathrm{A}},{\mathrm{B}})\simeq\lim_{{\mathrm{n}}\geq 1}\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})({\mathcal{O}}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{O}}_{\leq{\mathrm{n}}})}{\mathrm{A}},{\mathrm{B}})$
$\simeq\lim_{{\mathrm{n}}\geq 1}\mathrm{Alg}_{\mathsf{f}_{\mathrm{n}}({\mathcal{O}}_{\leq{\mathrm{n}}})}({\mathcal{C}})({\mathrm{A}},{\mathrm{B}})\simeq(\lim_{{\mathrm{n}}\geq 1}\mathrm{Alg}_{\mathsf{f}_{\mathrm{n}}({\mathcal{O}}_{\leq{\mathrm{n}}})}({\mathcal{C}}))({\mathrm{A}},{\mathrm{B}})$
@@
$\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\simeq\mathrm{Alg}_{\mathrm{colim}_{{\mathrm{n}}\geq 1}\mathsf{f}_{\mathrm{n}}({\mathcal{O}}_{\leq{\mathrm{n}}})}({\mathcal{C}})\simeq\lim_{{\mathrm{n}}\geq 1}\mathrm{Alg}_{\mathsf{f}_{\mathrm{n}}({\mathcal{O}}_{\leq{\mathrm{n}}})}({\mathcal{C}}).$
@@
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category compatible with the initial object that admits small colimits and let ${\mathrm{n}}\in\Sigma.$ Let ${\mathrm{X}}_{1},...,{\mathrm{X}}_{\mathrm{k}}$ be objects of ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ for ${\mathrm{k}}\geq 2.$ If ${\mathrm{X}}_{\mathrm{i}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}$ for some $1\leq{\mathrm{i}}\leq{\mathrm{k}},$ then ${\mathrm{X}}_{1}\circ...\circ{\mathrm{X}}_{\mathrm{k}}$ belongs to ${\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}$.
@@
Proposition 7.9 states that ${\mathcal{C}}$ embeds symmetric monoidally into a symmetric monoidal $\infty$-category ${\mathcal{C}}^{\prime}$ compatible with small colimits such that the embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ preserves the initial object and admits a left adjoint ${\mathrm{L}}.$ We obtain an embedding ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}_{\geq 1}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}_{\geq 1}$ of non-symmetric $\infty$-operads and an embedding ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{\mathrm{n}}}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}^{\leq{\mathrm{n}}}.$ Moreover ${\mathrm{X}}_{1}\circ...\circ{\mathrm{X}}_{\mathrm{k}}$ is the image of the corresponding composition product in ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}$ under the functor ${\mathrm{L}}_{\ast}:{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ induced by ${\mathrm{L}}.$ As a left adjoint the functor ${\mathrm{L}}$ preserves the initial object so that the functor ${\mathrm{L}}_{\ast}$ restricts to a functor ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq{\mathrm{n}}}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}$. Consequently, we can reduce to the case that the symmetric monoidal structure on ${\mathcal{C}}$ is compatible with small colimits. Let ${\mathrm{X}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$ For every $\mathsf{s}\geq 0$ we have
$({\mathrm{X}}\circ{\mathrm{Y}})_{\mathsf{s}}=\underset{{\mathrm{k}}\geq 0}{\coprod}(\underset{{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}=\mathsf{s}}}{\coprod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}=\underset{\mathsf{s}\geq{\mathrm{k}}\geq 0}{\coprod}(\underset{{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}=\mathsf{s}}}{\coprod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}$
so that ${\mathrm{X}}\circ{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}.$ Let ${\mathrm{X}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}.$ For every $\mathsf{s}\geq 0$ we have
$({\mathrm{X}}\circ{\mathrm{Y}})_{\mathsf{s}}=\underset{{\mathrm{k}}\geq 0}{\coprod}(\underset{{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}=\mathsf{s}}}{\coprod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}=\underset{{\mathrm{k}}\geq 1}{\coprod}(\underset{{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}=\mathsf{s}}}{\coprod}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}.$
Thus ${\mathrm{X}}\circ{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}.$
@@
For every ${\mathrm{n}}\geq 1$ let $\Sigma_{\geq{\mathrm{n}}}\subset\Sigma_{\geq 1}$ be the full subcategory of sets having at least ${\mathrm{n}}$ elements.

### v2

@@ 2.4. Truncating $\infty$-operads
For every ${n}\geq 1$ let $\Sigma_{\geq 1}^{\leq{n}}\subset\Sigma_{\geq 1}$ be the full subcategory of sets with less or equal than n{n} elements.
The embedding $\iota:\Sigma_{\geq 1}^{\leq{n}}\subset\Sigma_{\geq 1}$ induces a localization
$\iota^{\ast}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\rightleftarrows{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{n}}:\iota_{\ast},$
where the left adjoint is restriction and the fully faithful right adjoint extends beyond degree n{n} by the zero object.
@@
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category compatible with the zero object that admits small colimits. For every ${n}\geq 1$ the localization $\iota^{\ast}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\rightleftarrows{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{n}}:\iota_{\ast}$ is compatible with the composition product on ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$
@@
By Proposition 5.9 we find that $\mathcal{C}$} embeds symmetric monoidally into a symmetric monoidal $\infty$-category $\mathcal{C}$$^{\prime}$ compatible with small colimits such that the embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ preserves the zero object and admits a left adjoint $\mathrm{L}$}. We obtain an embedding ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}_{\geq 1}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}_{\geq 1}$ of non-symmetric $\infty$-operads and an embedding ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{n}}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}^{\leq{n}}.$ Thus ${\mathrm{X}}_{1}\circ...\circ{\mathrm{X}}_{k}$ is the image of the corresponding composition product in ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}$ under the functor ${\mathrm{L}}_{\ast}:{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ induced by $\mathrm{L}$}. A morphism of ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ is a local equivalence if and only if it induces an equivalence in degrees $\leq{n}$ and similar for $\mathcal{C}$$^{\prime}$. So the embedding ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}$ and its left adjoint ${\mathrm{L}}_{\ast}:{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ preserve local equivalences. Hence we can reduce to the case that the symmetric monoidal structure on $\mathcal{C}$} is compatible with small colimits.
Let ${\mathrm{f}}:{\mathrm{X}}\to{\mathrm{Y}}$ be a morphism in ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ such that ${\mathrm{f}}_{j}:{\mathrm{X}}_{j}\to{\mathrm{Y}}_{j}$ is an equivalence for every ${j}\leq{n}$ and ${\mathrm{Z}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$ We want to see that $({\mathrm{f}}\circ{\mathrm{Z}})_{s}$ and $({\mathrm{Z}}\circ{\mathrm{f}})_{s}$ are both equivalences for every $s\leq{n}.$ We compute
$({\mathrm{f}}\circ{\mathrm{Z}})_{s}=\underset{{k}\geq 0}{\coprod}(\underset{{{n}_{1}\coprod...\coprod{n}_{k}=s}}{\coprod}{\mathrm{f}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Z}}_{{n}_{j}}))_{\Sigma_{k}}=\underset{s\geq{k}\geq 0}{\coprod}(\underset{{{n}_{1}\coprod...\coprod{n}_{k}=s}}{\coprod}{\mathrm{f}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Z}}_{{n}_{j}}))_{\Sigma_{k}},$
$({\mathrm{Z}}\circ{\mathrm{f}})_{s}=\underset{{k}\geq 0}{\coprod}(\underset{{{n}_{1}\coprod...\coprod{n}_{k}=s}}{\coprod}{\mathrm{Z}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{f}}_{{n}_{j}}))_{\Sigma_{k}}.$
@@
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category that admits small colimits and a zero object preserved by the tensor product in each component. For every ${n}\geq 1$ let ${\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq{n}}\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}$ be the full subcategory of $\infty$-operads in $\mathcal{C}$} whose underlying symmetric sequence vanishes beyond degree n,{n}, i.e. belongs to the full subcategory ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{n}}\subset{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$
@@
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category that admits small colimits and a zero object preserved by the tensor product in each component. The embedding ${\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq{n}}\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}$ admits a left adjoint $(-)_{\leq{n}}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq{n}}$ that fits into a commutative square, where the bottom functor is restriction:
${{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{n}}.}$
@@
If $\mathcal{C}$} is a presentably symmetric monoidal $\infty$-category, by the adjoint functor theorem the functor $(-)_{\leq{n}}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq{n}}$ preserving small limits and sifted colimits admits a left adjoint denoted by ${\mathrm{f}}_{n}$. The composition
${\mathrm{f}}_{n}\circ(-)_{\leq{n}}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq{n}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}$
is left adjoint to the functor $\tau_{n}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\xrightarrow{(-)_{\leq{n}}}{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq{n}}\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}.$
@@
${\mathrm{f}}_{1}({\mathcal{O}}_{\leq 1})\to...\to{\mathrm{f}}_{n}({\mathcal{O}}_{\leq{n}})\to{\mathrm{f}}_{{n}+1}({\mathcal{O}}_{\leq{n}+1})\to...\to{\mathcal{O}}$
is a colimit diagram (see [30, Remark 4.23.] for details).
@@
${\mathcal{O}}\circ_{{\mathrm{f}}_{1}({\mathcal{O}}_{\leq 1})}{\mathrm{A}}\to...\to{\mathcal{O}}\circ_{{\mathrm{f}}_{n}({\mathcal{O}}_{\leq{n}})}{\mathrm{A}}\to{\mathcal{O}}\circ_{{\mathrm{f}}_{{n}+1}({\mathcal{O}}_{\leq{n}+1})}{\mathrm{A}}\to...\to{\mathcal{O}}\circ_{\mathcal{O}}{\mathrm{A}}\simeq{\mathrm{A}}$
@@
$\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})({\mathrm{A}},{\mathrm{B}})\to\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})(\mathrm{colim}_{{n}\geq 1}{\mathcal{O}}\circ_{{\mathrm{f}}_{n}({\mathcal{O}}_{\leq{n}})}{\mathrm{A}},{\mathrm{B}})\simeq\lim_{{n}\geq 1}\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})({\mathcal{O}}\circ_{{\mathrm{f}}_{n}({\mathcal{O}}_{\leq{n}})}{\mathrm{A}},{\mathrm{B}})$
$\simeq\lim_{{n}\geq 1}\mathrm{Alg}_{{\mathrm{f}}_{n}({\mathcal{O}}_{\leq{n}})}({\mathcal{C}})({\mathrm{A}},{\mathrm{B}})\simeq(\lim_{{n}\geq 1}\mathrm{Alg}_{{\mathrm{f}}_{n}({\mathcal{O}}_{\leq{n}})}({\mathcal{C}}))({\mathrm{A}},{\mathrm{B}})$
@@
$\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\simeq\mathrm{Alg}_{\mathrm{colim}_{{n}\geq 1}{\mathrm{f}}_{n}({\mathcal{O}}_{\leq{n}})}({\mathcal{C}})\simeq\lim_{{n}\geq 1}\mathrm{Alg}_{{\mathrm{f}}_{n}({\mathcal{O}}_{\leq{n}})}({\mathcal{C}}).$
@@
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category compatible with the initial object that admits small colimits and let ${n}\in\Sigma.$ Let ${\mathrm{X}}_{1},...,{\mathrm{X}}_{k}$ be objects of ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ for ${k}\geq 2.$ If ${\mathrm{X}}_{i}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}$ for some $1\leq{i}\leq{k},$ then ${\mathrm{X}}_{1}\circ...\circ{\mathrm{X}}_{k}$ belongs to ${\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}$.
@@
Proposition 5.9 states that $\mathcal{C}$} embeds symmetric monoidally into a symmetric monoidal $\infty$-category $\mathcal{C}$$^{\prime}$ compatible with small colimits such that the embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ preserves the initial object and admits a left adjoint $\mathrm{L}$}. We obtain an embedding ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}_{\geq 1}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}_{\geq 1}$ of non-symmetric $\infty$-operads and an embedding ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}^{\leq{n}}\subset{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}^{\leq{n}}.$ Moreover ${\mathrm{X}}_{1}\circ...\circ{\mathrm{X}}_{k}$ is the image of the corresponding composition product in ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}$ under the functor ${\mathrm{L}}_{\ast}:{\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ induced by $\mathrm{L}$}. As a left adjoint the functor $\mathrm{L}$} preserves the initial object so that the functor $\mathrm{L}$$_{\ast}$ restricts to a functor ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq{n}}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}$. Consequently, we can reduce to the case that the symmetric monoidal structure on $\mathcal{C}$} is compatible with small colimits. Let ${\mathrm{X}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$ For every $s\geq 0$ we have
$({\mathrm{X}}\circ{\mathrm{Y}})_{s}=\underset{{k}\geq 0}{\coprod}(\underset{{{n}_{1}\coprod...\coprod{n}_{k}=s}}{\coprod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))_{\Sigma_{k}}=\underset{s\geq{k}\geq 0}{\coprod}(\underset{{{n}_{1}\coprod...\coprod{n}_{k}=s}}{\coprod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))_{\Sigma_{k}}$
so that ${\mathrm{X}}\circ{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}.$ Let ${\mathrm{X}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}.$ For every $s\geq 0$ we have
$({\mathrm{X}}\circ{\mathrm{Y}})_{s}=\underset{{k}\geq 0}{\coprod}(\underset{{{n}_{1}\coprod...\coprod{n}_{k}=s}}{\coprod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))_{\Sigma_{k}}=\underset{{k}\geq 1}{\coprod}(\underset{{{n}_{1}\coprod...\coprod{n}_{k}=s}}{\coprod}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))_{\Sigma_{k}}.$
Thus ${\mathrm{X}}\circ{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}.$
@@
For every ${n}\geq 1$ let $\Sigma_{\geq{n}}\subset\Sigma_{\geq 1}$ be the full subcategory of sets having at least n{n} elements.